OpenStudy (anonymous):

prove that sin54-sin18=1/2

OpenStudy (goldrush18):

0.80902-0.30902=0.5 or 1/2

OpenStudy (goldrush18):

or should i say 0.5=1/2

OpenStudy (anonymous):

@GoldRush18 u have to prove it using trigonometric functions and rules , not using calculator :D

OpenStudy (goldrush18):

oh ok

OpenStudy (anonymous):

Use cofunction identity.

OpenStudy (anonymous):

well i got cos36-sin18 and dont know what to do , can u prove it with explanation pls :)

OpenStudy (anonymous):

Idk if this helps, but Sin(3theta)=3Sin(theta)-4(Sin(theta))^3 I tried using it and I got - 4 sin³ 18 + 2 sin 18 - 1/2 = 0

OpenStudy (blockcolder):

Use this identity: \[\sin A-\sin B=2 \cos{\frac{A+B}{2}}\sin{\frac{A-B}{2}}\]

OpenStudy (anonymous):

thanks will do that :D @blockcolder

OpenStudy (anonymous):

thanks i got it now it should be like this 2cos(36)sin(18) and multiply by cos18/cos18 as it equals 1 so cos(36)sin(36)/cos(18) and then multiply by 2 and 1/2 so 1/2sin72/cos18 then use co function so sin72= cos18 1/2cos(18)/cos(18) cosines cancel each other so = 1/2

OpenStudy (blockcolder):

Good job! :)